ALGEBRA 1
write in standard from: y+2=4/3(x+6)

Answer by rfer(16322)

y+2=4/3(x+6)
y+2=4/3x+8
subtract 2
y=4/3x+6
-4/3x+y=6
4/3x-y=-6